# Geometric Bounds for Persistence

### v1

@@ Appendix C Notation Table
$\mathrm{conv}(\mathcal{X})$ Convex hull of $\mathcal{X}\subset\mathbf{Y}$
$\overline{\mathrm{conv}}(\mathcal{X})$ Closure of convex hull of $\mathcal{X}\subset\mathbf{Y}$
@@
$\operatorname{\tilde{H}}_{k}(\cdot)$ Reduced singular homology in degree $k$
$\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$ Open $r$-neighborhood of $\mathcal{X}$ in $\mathbf{Y}$ Section 2.1
$\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathbf{Y})$ Closed $r$-neighborhood of $\mathcal{X}$ in $\mathbf{Y}$ Section 2.1
$\mathcal{V}_{r}(\mathcal{X})$ Vietoris-Rips complex of $\mathcal{X}$ for distance $r$ 2.1
$\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ech complex of $\mathcal{X}\subset\mathbf{Y}$ for distance $r$ 2.2
$\mathrm{Spec}_{k}(\Delta_{\bullet})$ Homological spectrum of the filtration $\Delta_{\bullet}$ 2.17
$\mathrm{PD}_{k}(\mathcal{X})$ $k^{th}$-persistence diagram of $\mathcal{X}$ with VR filtration Section 2.1.2
$\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ $k^{th}$-persistence diagram of $\mathcal{X}\subset\mathbf{Y}$ with Čech filtration Section 2.1.2
$b_{\omega},d_{\omega}$ Birth and death time of the homology class $\omega$ 2.18
$d_{\omega}-b_{\omega}$ Persistence (lifespan) of the homology class $\omega$ 2.18
@@
$\rho(\mathcal{M})$ Gromov’s filling radius of $\mathcal{M}$ 2.30
$\rho(\omega;\mathcal{X}\subset\mathbf{Y})$ The relative filling radius of $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$ relative to $\mathbf{Y}$ 2.31
$\rho(\omega;\mathcal{X})$ The absolute filling radius of $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$ 2.31
$\mathrm{rad}(\mathcal{X})$ The radius of $\mathcal{X}$ 2.35
$\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$ The circumradius of $\mathcal{X}$ relative to $\mathbf{Y}$ 2.35
@@
$\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$ $C$-robust $k$-dimensional treewidth of $\mathcal{X}$ 3.19

### v2

@@ Appendix C Notation Table
$\mathrm{conv}(\mathcal{X})$ Convex hull of $\mathcal{X}$ in a Banach space
$\overline{\mathrm{conv}}(\mathcal{X})$ Closure of convex hull of $\mathcal{X}$ in a Banach space
@@
$\widetilde{\mathrm{H}}_{k}(\cdot)$ Reduced singular homology in degree $k$
$\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})$ Open $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$ Section 2.1
$\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathcal{Z})$ Closed $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$ Section 2.1
$\mathcal{V}_{r}(\mathcal{X})$ Vietoris–Rips complex of $\mathcal{X}$ for distance $r$ 2.1
$\mathcal{C}_{r}(\mathcal{X}\subset\mathcal{Z})$ech complex of $\mathcal{X}\subset\mathcal{Z}$ for distance $r$ 2.2
$\mathrm{Spec}_{k}(\Delta_{\bullet})$ Homological spectrum of the filtration $\Delta_{\bullet}$ 2.18
[…]
$b_{\omega},d_{\omega}$ Birth and death time of the homology class $\omega$ 2.19
$d_{\omega}-b_{\omega}$ Persistence (lifespan) of the homology class $\omega$ 2.19
@@
$\operatorname{E}(\mathcal{X})$ Tight span of $\mathcal{X}$ 2.28
$\rho(\mathcal{M})$ Gromov’s filling radius of $\mathcal{M}$ 2.31
$\rho(\omega;\mathcal{X}\subset\mathbf{Y})$ The relative filling radius of $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$ relative to $\mathbf{Y}$ 2.33
$\rho(\omega;\mathcal{X})$ The absolute filling radius of $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$ 2.33
$\mathrm{rad}(\mathcal{X})$ The radius of $\mathcal{X}$ 2.39
$\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$ The circumradius of $\mathcal{X}$ relative to $\mathbf{Y}$ 2.39
@@
$\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$ $C$-robust $k$-dimensional treewidth of $\mathcal{X}$ 3.20